If 15x is 20, what is 10x
Found 2 solutions by josgarithmetic, Alan3354:
Answer by josgarithmetic(39616)
Answer by Alan3354(69443)

If 15x is 20, what is 10x
-----------
You can solve for x, then do the rest.
15x = 20
x = 20/15 = 4/3
---
--> 10x = 40/3
===================
Or,
10x = 2/3 of 15x = 40/3